# Sculptor: Empowering LLMs with Cognitive Agency via Active Context Management

### v1

@@ Abstract
Large Language Models (LLMs) suffer from significant performance degradation when processing long contexts due to proactive interference, where irrelevant information in earlier parts of the context disrupts reasoning and memory recall. While most research focuses on external memory systems to augment LLMs’ capabilities, we propose a complementary approach: empowering LLMs with Active Context Management (ACM) tools to actively sculpt their internal working memory. We introduce Sculptor, a framework that equips LLMs with three categories of tools: (1) context fragmentation, (2) summary, hide, and restore, and (3) intelligent search. Our approach enables LLMs to proactively manage their attention and working memory, analogous to how humans selectively focus on relevant information while filtering out distractions. Experimental evaluation on information-sparse benchmarks—PI-LLM (proactive interference) and NeedleBench Multi-Needle Reasoning—demonstrates that Sculptor significantly improves performance even without specific training, leveraging LLMs’ inherent tool calling generalization capabilities. By enabling Active Context Management, Sculptor not only mitigates proactive interference but also provides a cognitive foundation for more reliable reasoning across diverse long-context tasks—highlighting that explicit context-control strategies, rather than merely larger token windows, are key to robustness at scale.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across diverse tasks, yet they face fundamental challenges when processing long contexts. Prior work shows that simply enlarging the context window leaves models vulnerable to position bias, overload, and interference as sequences grow (Liu et al., 2023; Hsieh et al., 2024a). Recent studies (Wang and Sun, 2025) have empirically demonstrated that LLMs suffer from proactive interference, where earlier information in the context disrupts the processing of subsequent, more relevant information. Moreover, calibrations like Found in the Middle (Hsieh et al., 2024b) reduce—but do not eliminate—positional bias; recent evaluations (Tian et al., 2025) find that performance still degrades significantly when the distance between relevant information pieces increases, as irrelevant information between them interferes with effective information integration. These phenomena mirror human cognitive psychology, where new learning can be impaired by previously acquired information that is no longer relevant to the current task.
The challenge becomes particularly acute in complex, multi-step reasoning tasks where LLMs must maintain focus on multiple critical information pieces while filtering out contextual noise (Li et al., 2025a). Traditional approaches to address long-context challenges have primarily focused on expanding context windows or developing external memory systems (Li et al., 2025c; Wang and Chen, 2025; Chhikara et al., 2025). While these solutions increase the amount of information an LLM can access, they do not address the fundamental issue of proactive interference—the inability to actively manage and curate the working memory that directly influences reasoning processes.
Consider a human expert working on a complex problem: they naturally employ active memory management strategies, selectively attending to relevant information, summarizing key insights, and temporarily setting aside less important details. They can revisit previously discarded information when needed, but crucially, they do not allow irrelevant details to continuously interfere with their current reasoning process. Current LLMs lack this fundamental cognitive capability. We propose that the solution lies not merely in expanding memory capacity, but in empowering LLMs with the ability to actively manage their internal working memory. Unlike external memory systems that store information outside the model’s immediate context, we focus on optimizing the model’s working memory—the immediate working space where attention operates and reasoning occurs.
@@
Figure 1: Overview of Sculptor framework: LLMs transform from being overwhelmed by cluttered context (left) to successfully solving solving tasks within noisy contexts by actively curating their working memory through context folding and organization (right).
This approach represents a paradigm shift from passively processing ever-growing contexts to active context curation. Instead of being overwhelmed by increasingly long contexts, LLMs learn to proactively manage their attention and working memory, focusing computational resources on the most relevant information. We view Sculptor as a representative of this emerging direction—complementary to external memory and context extension—providing a necessary step toward reliable long-horizon reasoning. Related work on context compression (Xu et al., 2023; Jiang et al., 2024) further demonstrates that selectively foregrounding key information can simultaneously improve accuracy and reduce cost/latency, reinforcing the need for explicit context control over passive attention alone. Recent work also suggests that in-context learning can be viewed as implicit weight updates (Dherin et al., 2025), implying that allowing models to modify their own context enables a form of “self-evolution” (Zhang et al., 2025)—a step toward agents that can adapt their computational substrate without external intervention.
@@
  • We conduct preliminary evaluations on two information-sparse benchmarks: PI-LLM for proactive interference and NeedleBench Multi-Needle Reasoning for information retrieval and reasoning, demonstrating promising improvements over baseline approaches.
  • We discuss ACM limitations, particularly the computational trade-offs from context reshaping, and outline future work including RL-based approaches to enhance tool capabilities.
2 Related Work

### v2

@@ Abstract
Large Language Models (LLMs) suffer from significant performance degradation when processing long contexts due to proactive interference, where irrelevant information in earlier parts of the context disrupts reasoning and memory recall. While most research focuses on external memory systems to augment LLMs’ capabilities, we propose a complementary approach: empowering LLMs with Active Context Management (ACM) tools to actively sculpt their internal working memory. We introduce Sculptor, a framework that equips LLMs with three categories of tools: (1) context fragmentation, (2) summary, hide, and restore, and (3) precise search. Our approach enables LLMs to proactively manage their attention and working memory, analogous to how humans selectively focus on relevant information while filtering out distractions. Experimental evaluation on diverse long-context benchmarks demonstrates that Sculptor significantly improves performance even without specific training, leveraging LLMs’ inherent tool-calling and instruction-following capabilities. To further optimize these strategies, we introduce a novel dynamic context-aware reinforcement learning (RL) approach, advancing the training of an agent that actively modifies its own conversational history. By enabling Active Context Management, Sculptor not only mitigates proactive interference but also provides a cognitive foundation for more reliable reasoning across diverse long-context tasks—highlighting that explicit context-control strategies, rather than merely larger token windows, are key to robustness at scale.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across diverse tasks, yet they face fundamental challenges when processing long contexts. Prior work shows that simply enlarging the context window leaves models vulnerable to position bias, overload, and interference as sequences grow (Liu et al., 2023a; Hsieh et al., 2024a). Recent studies (Wang & Sun, 2025) have empirically demonstrated that LLMs suffer from proactive interference, where earlier information in the context disrupts the processing of subsequent, more relevant information. Moreover, calibrations like Found in the Middle (Hsieh et al., 2024b) reduce—but do not eliminate—positional bias; recent evaluations (Tian et al., 2025) find that performance still degrades significantly when the distance between relevant information pieces increases, as irrelevant information between them interferes with effective information integration. These phenomena mirror human cognitive psychology, where new learning can be impaired by previously acquired information that is no longer relevant to the current task.
The challenge becomes particularly acute in complex, multi-step reasoning tasks where LLMs must maintain focus on multiple critical information pieces while filtering out contextual noise (Li et al., 2025a). Traditional approaches to address long-context challenges have primarily focused on expanding context windows or developing external memory systems (Li et al., 2025c; Wang & Chen, 2025; Chhikara et al., 2025). While these solutions increase the amount of information an LLM can access, they do not address the fundamental issue of proactive interference—the inability to actively manage and curate the working memory that directly influences reasoning processes.
Consider a human expert working on a complex problem: they naturally employ active memory management strategies, selectively attending to relevant information, summarizing key insights, and temporarily setting aside less important details. They can revisit previously discarded information when needed, but crucially, they do not allow irrelevant details to continuously interfere with their current reasoning process. Current LLMs lack this fundamental cognitive capability. We propose that the solution lies not merely in expanding context window, but in empowering LLMs with the ability to actively manage their internal working memory. Unlike external memory systems that store information outside the model’s immediate context, we focus on optimizing the model’s working memory—the immediate working space where attention operates and reasoning occurs.
@@
Figure 1: Overview of Sculptor framework: Through Active Context Management, LLMs transform overloaded contexts into optimized contexts using fragment, summary, search, and restore operations, enabling successful task completion where traditional approaches fail due to interference.
This approach represents a paradigm shift from passively processing ever-growing contexts to active context curation. Instead of being overwhelmed by increasingly long contexts, LLMs learn to proactively manage their attention and working memory, focusing computational resources on the most relevant information. We view Sculptor as a representative of this emerging direction—complementary to external memory and context extension—providing a necessary step toward reliable long-horizon reasoning. Related work on context compression (Xu et al., 2023; Jiang et al., 2024b; Guo et al., 2025) further demonstrates that selectively foregrounding key information can simultaneously improve accuracy and reduce cost and latency, reinforcing the need for explicit context control over passive attention alone. Recent work also suggests that in-context learning can be viewed as implicit weight updates (Dherin et al., 2025), implying that allowing models to modify their own context enables a form of “self-evolution” (Zhang et al., 2025)—a step toward agents that can adapt their computational substrate without external intervention.
@@
  • We propose an RL training approach for active context modification, introducing Conditional Trajectory Collection and Incremental Loss Assignment to enable effective learning of context manipulation strategies. Through dynamic context-aware GSPO training, we achieve substantial performance gains across diverse long-context benchmarks.
  • We provide comprehensive analysis of tool usage patterns, attention mechanisms, and cost analysis, demonstrating that ACM effectively reduces context token consumption while enhancing long-context capabilities.
2 Methodology
2.1 Tool Design Principles
Our tool design follows four core principles. (1) Deterministic and Self-Contained Operations: each tool is a simple, deterministic operator without external dependencies (e.g., embedding models), a self-contained design that guarantees deployment stability and isolates the LLM’s cognitive agency for pure evaluation. (2) Cognitive Alignment: the tools mirror effective human strategies, such as our search_context tool performing exact matching akin to “Ctrl+F”, a computationally efficient approach that reserves complex semantic understanding for the LLM’s own reasoning. (3) Structural Preservation for Scalable Training: the tools are constrained to never alter the count or order of messages, thereby maintaining a stable state representation that is critical for tractable credit assignment in reinforcement learning. (4) Reversibility and Graceful Degradation: All context-modifying operations are designed to be non-destructive and fully reversible (e.g., fold is undone by expand), ensuring no information is permanently lost. This guarantees that the framework functions as a strict superset of the baseline model’s capabilities, allowing for graceful degradation: if no tools are invoked, the model’s behavior is identical to its original, unmodified state.
2.2 The Sculptor Tool Suite
Following these design principles, we equip LLMs with six fundamental tools organized into three functional categories, allowing them to work in coordination within a single turn, where the agent receives a user message and performs multi-step tool calls—for instance, fragmenting a context segment yields a unique fragment ID that enables subsequent operations like compression, summarization, or restoration—continuously invoking these tools until generating a final response without further tool invocations. Complete JSON schemas for all tools are provided in Appendix F.
(2) Context Compression and Restoration involves three complementary tools for dynamic content management. summarize_fragment generates focused AI-powered summaries of specific fragments based on user-specified focus areas (e.g., technical details, key decisions, action items), compressing content while preserving critical information. fold_fragment temporarily hides fragment content while preserving its existence, displaying only a folded marker to dramatically reduce visual clutter. restore_fragment provides universal restoration capability, reverting both summarized and folded fragments back to their original content, ensuring no information is permanently lost during context management operations.
(3) Precise Search and Retrieval is accomplished through two complementary tools. search_context performs exact keyword matching across user messages, assistant responses, or all content—mirroring the human approach of using Ctrl+F for information retrieval. It returns up to 50 matches with configurable result context windows. get_search_detail retrieves extended context around specific search results, with the model specifying the desired surrounding character count. By appending search results to the end of conversation history, this approach mitigates the “lost in the middle” problem (Liu et al., 2023a) where models struggle to locate information buried within long contexts.
3 Teaching LLMs to Use Sculptor Tools
Building on the strong tool-use capabilities inherent in modern LLMs, we explore two distinct approaches for teaching models to effectively wield the Sculptor tool suite. Throughout this paper, we use “ACM tools” to specifically refer to our Sculptor implementation—a concrete instantiation of the broader Active Context Management paradigm.
3.1 Inherent Tool-Use Performance
We first evaluate the inherent tool-calling capabilities of state-of-the-art models like Claude-4-Sonnet and GPT-4.1, which demonstrate strong zero-shot generalization abilities for function calling. These models can understand and execute our Sculptor tools without any specific training, relying on their pre-trained understanding of tool usage patterns and natural language descriptions of tool schema. This zero-shot approach requires no additional training—models directly interpret and use the tools based solely on their schemas. To encourage consistent tool engagement, we set tool_choice=“required” for the first step of multi-step conversations.
3.2 Multi-step Agent RL Training with Dynamic Context-Aware GSPO
To optimize tool usage strategies beyond zero-shot generalization, we develop a reinforcement learning approach specifically designed for multi-step tool calling in long-context scenarios. Our approach addresses the unique challenges of training models to actively manage dynamic contexts where tool calls can fundamentally alter the information landscape.
Group Sequence Policy Optimization (GSPO).
We adapt GSPO (Zheng et al., 2025) for multi-step rl training, leveraging its sequence-level optimization for stable training in long-context scenarios. Given a query $x$ and $G$ sampled trajectories $\{\tau_{i}\}_{i=1}^{G}$ from policy $\pi_{\theta_{\text{old}}}$, GSPO optimizes:
$\mathcal{J}_{\text{GSPO}}(\theta)=\mathbb{E}_{x\sim\mathcal{D}}\left[\frac{1}{G}\sum_{i=1}^{G}\min\left(s_{i}(\theta)\hat{A}_{i},\text{clip}(s_{i}(\theta),1-\varepsilon,1+\varepsilon)\hat{A}_{i}\right)\right]$ (1)
where the sequence-level importance ratio is:
$s_{i}(\theta)=\left(\frac{\pi_{\theta}(\tau_{i}|x)}{\pi_{\theta_{\text{old}}}(\tau_{i}|x)}\right)^{\frac{1}{|\tau_{i}|}}$ (2)
and the group-normalized advantage is:
$\hat{A}_{i}=\frac{r(x,\tau_{i})-\text{mean}(\{r(x,\tau_{j})\}_{j=1}^{G})}{\text{std}(\{r(x,\tau_{j})\}_{j=1}^{G})}$ (3)
Dynamic Context-Aware Credit Assignment with Incremental Loss Design.
The key innovation in our approach addresses the non-monotonic nature of context evolution during tool calling. Traditional multi-step RL assumes each trajectory $\tau_{t}$ is a prefix of $\tau_{t+1}$, allowing training only on the final trajectory. However, with context management tools, $c_{t}\not\subset c_{t+1}$ in general—tools like fold_fragment or summarize_fragment actively remove or transform information, creating divergent context states.
To handle this, we introduce a two-part strategy combining conditional trajectory collection and incremental loss assignment, illustrated in Figure 2 and detailed in Appendix D. The final reward is propagated to all sub-trajectories within the same rollout, ensuring each context state receives appropriate learning signal. This incremental design prevents the model from learning spurious patterns where context-modifying tools repeatedly trigger themselves, which would cause training collapse. Each tool call receives gradient signal exactly once across all completions, ensuring stable and efficient learning. Notably, this method applies equally to both supervised fine-tuning (SFT) and reinforcement learning stages, providing a unified framework for training with dynamic contexts.
Figure 2: Conditional trajectory collection and incremental loss assignment for RL training. Q represents the initial user context, C denotes assistant completions, and T indicates tool results. Top: Complete interaction trajectory with context-modifying tool at step C2. Bottom: Training samples extracted via conditional trajectory collection, where each context change creates a new training instance. Incremental loss is assigned only to new completions (red boxes) while masking prior completions (loss=0), preventing redundant learning and training collapse.
4 Experiments
We evaluate Sculptor in two settings: zero-shot tool calling leveraging models’ inherent capabilities, and after reinforcement learning with dynamic context-aware GSPO to optimize tool usage strategies.
4.1 Evaluating Prompt-Guided Tool Calling Performance
Evaluated Models:
Evaluated Benchmarks:
We evaluate on five benchmarks testing diverse long-context challenges: (1) PI-LLM (Wang & Sun, 2025) tests proactive interference through continuous key-value updates (2-256 updates, 46 keys). (2) NeedleBench (Li et al., 2025a) Multi-Needle Reasoning requires connecting 2-5 needles simultaneously across varying context lengths. For cost efficiency and rapid validation, we initially evaluate only on PI-LLM and NeedleBench in zero-shot settings. After RL training, we expand to: (3) MRCR (Vodrahalli et al., 2024a) for multi-round co-reference resolution, requiring models to distinguish between multiple identical requests (2-8 needles) and return the i-th occurrence from synthetic conversations. (4) LongBenchV2 (Bai et al., 2025) for comprehensive long-context understanding. (5) FRAMES (Krishna et al., 2025) for factuality, retrieval, and reasoning measurement, containing 824 multi-hop questions requiring integration of information from 2-15 Wikipedia articles.
Inherent Challenges of Unguided Tool Use:
To understand how models naturally interact with ACM tools, we conducted initial experiments using Claude-4-Sonnet on PI-LLM and NeedleBench benchmarks, collecting 50 samples from each benchmark for tool usage analysis. We provided the model with a unified system prompt—minimal, generic instructions applicable across all tasks—without any benchmark-specific guidance (see Section C.2 for the complete prompt).
Figure 3: Tool usage count comparison for Claude-4-Sonnet before and after prompt optimization. Without task-specific prompts (unified prompt), both benchmarks show suboptimal patterns. With benchmark-specific prompt engineering, distinct improvements emerge: PI-LLM shifts from inefficient search-heavy patterns (625 calls) to strategic fold_fragment usage (1206 calls) for managing obsolete information, while NeedleBench increases search operations from 77 to 206 calls—addressing insufficient execution depth through more thorough verification and multi-hop reasoning. These contrasting patterns highlight how prompt engineering resolves different challenges: tool selection efficiency for PI-LLM versus execution completeness for NeedleBench.
Our findings revealed suboptimal tool selection patterns, as shown in Figure 3 (left bars). For PI-LLM, which contains numerous obsolete key-value pairs requiring the model to focus on the latest mappings, we expected the model to leverage fragment_context and fold_fragment to compress outdated information. However, Claude-4-Sonnet overwhelmingly relied on search_context (90.7% of tool calls), attempting exhaustive searches for each of the 46 keys despite hundreds of historical updates per key. This search-heavy approach proved highly inefficient—the model exhausted its 20-tool-call budget merely aggregating occurrences without effectively filtering obsolete information. Similarly, for NeedleBench, while search tools are appropriate for retrieval tasks, the model showed limited strategic diversity in tool selection.
These observations reveal three fundamental challenges in unguided tool calling: (1) Suboptimal tool selection efficiency: The model failed to recognize when certain tools become inefficient for specific scenarios. In PI-LLM, attempting exhaustive searching for 46 keys with hundreds of historical updates each consumed the entire tool budget, when structural reorganization through fragment-and-fold would have been far more efficient. (2) Tool dependency misunderstanding: The model lacked comprehension of tool prerequisites and operational dependencies—for instance, attempting to use summary_by_id before generating fragment IDs with fragment_context, demonstrating incomplete understanding of the tool suite’s workflow. (3) Insufficient execution depth: Even when correctly initiating tool usage, the model often failed to complete tasks thoroughly, with incomplete fragmentation where only partial sections were processed, leaving critical information unaddressed. These challenges underscore that effective ACM tool usage requires not just access to tools but deep understanding of efficiency trade-offs, operational dependencies, and thorough execution strategies.
From Baseline Struggles to Systematic Guidance:
To address these inefficiencies, we crafted benchmark-specific prompts to steer tool strategies: for PI-LLM, first fragment then fold before any search or answering; for NeedleBench, coordinate search_context and get_search_detail for multi-hop retrieval. As shown in Figure 3, this guidance shifts patterns accordingly (PI-LLM: from search-heavy to fragment+fold; NeedleBench: deeper search), improving tool selection efficiency and execution completeness.
This systematic prompt engineering approach also enabled us to generate high-quality training data, collecting numerous successful tool usage examples across benchmarks. These guided patterns demonstrate that proper instruction can unlock more effective tool utilization, transforming suboptimal default behaviors into strategic, task-appropriate tool selection. The complete system prompt templates used in our experiments are provided in Sec. C.1.
Table 1: Performance improvements of frontier models with ACM Tools on NeedleBench-M-RS and PI-LLM benchmarks. Both benchmarks demonstrate substantial performance gains.
Method NeedleBench-M-RS PI-LLM (Update Count / Context Length)
2-N 3-N 4-N 5-N Avg 4/1K 8/2K 16/4K 32/8K 64/16K 128/32K 256/64K Avg
Claude-4-Sonnet
Baseline 96.0 82.0 54.0 36.0 67.0 99.13 95.65 92.17 84.78 81.74 65.22 69.57 84.04
w/ ACM Tools 100.0 98.0 88.0 90.0 94.0 90.43 91.74 98.26 92.17 91.74 87.39 77.83 89.94
$\Delta$ +4.0 +16.0 +34.0 +54.0 +27.0 -8.70 -3.91 +6.09 +7.39 +10.00 +22.17 +8.26 +5.90
GPT-4.1
Baseline 90.0 64.0 30.0 8.0 48.0 96.96 91.30 79.57 67.83 63.04 63.91 50.43 73.29
w/ ACM Tools 96.0 84.0 60.0 44.0 71.0 92.17 89.13 93.04 83.91 76.09 64.35 60.43 79.87
$\Delta$ +6.0 +20.0 +30.0 +36.0 +23.0 -4.79 -2.17 +13.47 +16.08 +13.05 +0.44 +10.00 +6.58
DeepSeek-V3
Baseline 88.0 68.0 28.0 16.0 50.0 95.22 85.65 70.00 63.91 33.04 32.17 21.74 57.39
w/ ACM Tools 92.0 58.0 50.0 32.0 58.0 73.91 90.00 79.13 37.39 53.04 55.65 11.74 57.27
$\Delta$ +4.0 -10.0 +22.0 +16.0 +8.0 -21.31 +4.35 +9.13 -26.52 +20.00 +23.48 -10.00 -0.12
Figure 4: NeedleBench Multi-Needle Reasoning performance across different context lengths. Left: Performance by needle count showing both with tool and vanilla results. Right: Average performance across all needle counts demonstrating significant improvements.
Performance Results:
Table 1 presents the evaluation results comparing models with and without ACM tools, using optimized benchmark-specific prompts. The improvements demonstrate the power of combining ACM tools with proper guidance: On NeedleBench-M-RS, Claude-4-Sonnet, GPT-4.1, and DeepSeek-V3 achieve gains of 27.0, 23.0, and 8.0 points respectively when using ACM tools with task-specific prompts, with Claude-4-Sonnet reaching 90% accuracy on 5-needle tasks. For PI-LLM, Claude-4-Sonnet and GPT-4.1 gain 5.90 and 6.58 points, while DeepSeek-V3 shows a slight decrease (-0.12), revealing persistent challenges even with prompt optimization. These results demonstrate that while prompt engineering significantly improves tool utilization, the degree of improvement varies based on each model’s inherent tool-use capabilities, suggesting the need for more systematic training approaches.
4.2 Optimizing Tool Use with Reinforcement Learning
While prompt engineering enables effective tool usage, it requires manual effort to design task-specific prompts and still exhibits the inherent limitations discussed above. To address these challenges systematically, we employ reinforcement learning to train models that can autonomously determine optimal tool usage strategies without explicit guidance.
Model and baselines.
We base our experiments on M3, a 13B-parameter dense model that we pre-train from scratch. We choose M3 for its strong tool-use capabilities (see Appendix B), tight compatibility with our training infrastructure, and competitive baseline performance. In Table 2, we additionally compare two baseline methods (both implemented on the same M3 base model for a controlled comparison): retrieval-augmented generation (RAG) and MemAgent(Yu et al., 2025). For fairness and to isolate the LLM’s inherent capabilities, RAG uses BM25-only retrieval without any external embedding models; Further evaluation details are provided in Sec. C.4.
Training Data Collection.
While M3 possesses strong inherent tool-use capabilities, it requires specific training to effectively utilize the Sculptor tools. We generate high-quality training data through the systematic prompt engineering approach described in Section 4.1. Using Claude-4-Sonnet with carefully designed task-specific prompts, we collect successful tool usage trajectories on the BABILong (Kuratov et al., 2024) and GSM-Infinite (Zhou et al., 2025) datasets—public benchmarks featuring complex long-context reasoning challenges. This process yields diverse examples of effective ACM tool usage patterns across different task types. Combined with our conditional trajectory collection and incremental loss assignment methodology (Section 3.2), we first perform supervised fine-tuning on this data to obtain Sculptor-M3, which has learned basic ACM tool capabilities. Subsequently, we conduct RL training with dynamic context-aware GSPO on the same datasets to obtain Sculptor-M3-RL, enabling the model to autonomously discover optimal tool usage strategies. During training, we cap tool steps at 20 per turn, matching Claude-4-Sonnet’s effective zero-shot usage while keeping rollouts efficient.
Table 2: Main experimental results across benchmarks. M3 indicates our 13B baseline model without ACM tools. Sculptor-M3 is equipped with Sculptor tools and fine-tuned on ACM-specific data. Sculptor-M3-RL is further trained with dynamic context-aware GSPO. Bold indicates best performance, underline indicates second best.
Method PI-LLM NeedleBench-M-RS MRCR LongBenchV2 Frames Avg
(Acc %) (Acc %) (Acc %) (Acc %) (Acc %) (Norm)
M3 (Baseline) 22.5 30.0 46.3 33.0 65.2 39.4
M3 + RAG 17.9 12.5 6.6 25.8 33.6 19.3
M3 + MemAgent 41.5 24.0 22.1 29.6 61.5 35.7
Sculptor-M3 71.8 67.6 79.1 29.2 51.2 59.8
Sculptor-M3-RL 99.4 84.8 85.7 34.5 64.6 73.8
RL Training Results:
Table 2 presents our experimental results. Sculptor-M3 shows improvements over baseline M3, particularly on PI-LLM (+49.3 points), NeedleBench-M-RS (+37.6 points), and MRCR (+32.8 points). After GSPO training on BABILong and GSM-Infinite datasets, Sculptor-M3-RL reaches 99.4% on PI-LLM with gains across most benchmarks.
The baseline methods show clear limitations: MemAgent achieves 41.5% on PI-LLM but only 24.0% on NeedleBench-M-RS, as its query-dependent memory accumulation discards information that appears irrelevant initially but proves critical for multi-hop reasoning. RAG performs even worse due to BM25’s reliance on direct keyword matching, missing indirectly related context. Both methods suffer from irreversible filtering decisions based solely on the final query. In contrast, our ACM tools enable reversible context management—folding currently irrelevant information while preserving restoration capability. This flexibility explains the strong performance on PI-LLM, NeedleBench-M-RS, and MRCR. The modest gains on LongBenchV2 (+1.5 points) and slight decrease on FRAMES (-0.6 points) suggest comprehensive reasoning tasks benefit less from selective attention.
4.3 Value-Specific Attention Analysis
Figure 5: Value-specific attention analysis results. Left: Scatter plot comparing attention weights between folded and complete versions for 46 key-value pairs. Most points lie above the equality line, indicating improved attention with folding. Right: Distribution of attention improvements, showing a clear positive shift and confirming the systematic benefit of our approach.
Figure 6: Average context length reduction with Sculptor across benchmarks. Arrows indicate reduction percentages achieved through strategic tool usage.
To precisely quantify how content folding impacts attention allocation to critical information, we conduct a token-level value-specific attention analysis. While traditional approaches assume that attention mechanisms naturally learn to ignore irrelevant information during pretraining, our analysis reveals that explicitly removing distracting content significantly enhances attention focus. The core idea is to measure the attention from the tokens of a specific critical value in the model’s response back to the corresponding tokens of the same value in the input context. Our experiment uses 46 predefined key-value pairs from the PI-LLM benchmark as the critical information. For each pair, we calculate the attention score by first identifying the exact token positions of the value in both the input and the response. We then aggregate the attention weights across all layers and heads, averaging them to produce a single score that represents the model’s focus on that specific piece of information. This allows for a direct comparison between the “folded” context and the “complete” context scenarios. The results presented in Figure 6 demonstrate a significant and systematic improvement in attention allocation. Out of 46 key-value pairs, 43 (93.5%) exhibited enhanced attention in the folded version, achieving a mean improvement of 9.87% (ranging from -6.86% to +37.56%). The scatter plot reveals a strong positive correlation (R² = 0.97) with the vast majority of data points positioned above the equality line, confirming that the improvements are consistent and not random.
Notable performance gains were observed for pairs such as “law: contract” (+37.56%), “climate: heat dome” (+30.44%), and “emotion: indifferent” (+25.83%). A one-sample t-test on the distribution of improvements confirms that they are statistically significant (p ¡ 0.001), with a median improvement of 9.9%. These findings provide strong empirical evidence that folding redundant content enhances attention allocation to critical information by reducing attention dilution—even in well-pretrained models, irrelevant information interferes with attention mechanisms rather than being naturally filtered out. The consistent improvements across diverse semantic categories suggest that explicit context management through folding is more effective than relying solely on learned attention patterns.
4.4 Cost Analysis
To evaluate the computational efficiency of our approach, we analyze the context reduction achieved by Sculptor-M3-RL on benchmarks containing substantial irrelevant information. As shown in Figure 6, Sculptor-M3-RL achieves dramatic context reductions across these challenging benchmarks: 76.2% reduction on NeedleBench-M-RS (from 85.2K to 20.3K tokens), 89.0% on PI-LLM (from 72.4K to 8.0K tokens), and 79.1% on MRCR (from 95.6K to 20.0K tokens). These substantial reductions directly translate to computational savings, as the quadratic complexity of attention mechanisms makes processing cost heavily dependent on context length.
Importantly, our tool design minimizes additional computational overhead. Read-only tools like search_context preserve the prefix relationship between completions and fully reuse KV cache—they only add a few search operations while most of the context remains cached. This is similar to traditional tool use where KV cache can be efficiently reused. For context-modifying tools that do break the prefix relationship, the dramatic context reduction itself compensates for the cache invalidation cost. Processing 20K tokens even without caching is significantly faster than processing 85K tokens with full caching. This design—separating context-preserving search tools from context-modifying compression tools—ensures that our system achieves substantial context reduction with minimal computational overhead.
5 Limitations and Future Work
Our study primarily targets long-context scenarios, but ACM is also promising beyond long contexts. In mathematical reasoning, early mistakes can cascade due to autoregressive “prefix lock-in” that degrades subsequent correctness; folding or suppressing erroneous early steps may reset the trajectory and improve robustness (Wang & Sun, 2025; Feng et al., 2025; Wen et al., 2025). Future work will extend ACM to non-long-context domains (e.g., math, coding) and pursue richer training strategies and reward design to learn finer-grained tool-use policies, with the goal of improving performance on complex long-context benchmarks where our current results remain modest, such as LongBenchV2 and FRAMES (Bai et al., 2025; Krishna et al., 2025).
Appendix A Related Work